OpenStudy (anonymous):

Solve. 7x^2 – 63 = 0 x = 7 and x = 9 x = 9 and x = 3 x = 3 and x = –3 x = 3 and x= 7

OpenStudy (anonymous):

x=3,-3

OpenStudy (anonymous):

7x^2 = 63 x^2 = 9 x = 3, -3

OpenStudy (angela210793):

7(x^2-9)=0 x^2-9=0 x^2=9 x=+ - 3